Compton scattering can occur from atoms as well as from electrons.

Part A

If photons scatter from atoms of helium (He) gas, what's their maximum wavelength change?

Express your answer with the appropriate units.
Answer #1

Mass:

Mass of the helum atom, m = 6.645 x 10^-27 kg

Compton scattering:

Change in wavelength, \Delta \lambda = ( h / m c ) [ 1 - cos \theta ]

For maximum wavelength change: \theta = 180 degrees

Maximum wavelength change:

\Delta \lambda =  { ( 6.63 x 10^-34 ) / [ ( 6.645 x 10^-27 ) * ( 3 x 10^8) ]  } * [ 1 - (-1) ]

        = 0.665 x 10^-15 m